### v1

@@ 2.5. Hodge bundles
Suppose that $(Y,D)$ is a projective nc log scheme over $k$ as above. Let $\overline{f}:(X,E)\to(Y,D)$ be a projective log smooth morphism to $(Y,D)$, which restricts to an abelian scheme $f:A\to Y-D$ over $Y-D$. Let $\overline{F}^{1}=\overline{f}_{*}\Omega^{1}_{(X,E)/(Y,D)}$.
@@
Without loss of generality, we can assume that $k={\mathbb{C}}$. Switching to the analytic topology, $R^{1}f^{an}_{*}{\mathbb{C}}$ underlies a polarized variation of Hodge structure $(V,F^{\bullet},\ldots)$ of type $\{(1,0),(0,1)\}$, with unipotent monodromy about components of $D$. The Hodge bundle $F^{1}V=f^{an}_{*}\Omega^{1}_{A/Y-D}$ extends to $(\overline{F}^{1})^{an}$ over $Y$. The theorem now follows from the much more general result due to Fujino-Fujisawa-Saito [12, theorem 3]. ∎
@@
2.6. Beilinson’s gluing construction
We recall a basic result of Beilinson [6, 24, 27]. We will follow the notation from [7] and write $f_{*}$ etc. for the derived direct image. Also, $\psi_{f},\phi_{f}$ denotes derived nearby/vanishing cycle functors shifted by $[-1]$ (so as to preserve perverse sheaves). Let $\psi^{u}_{f}\subset\psi_{f},\phi^{u}_{f}\subset\phi_{f}^{u}$ denote the maximal subfunctors on which the monodromy $T$ acts unipotently. The $can$ and $var$ maps restrict to these. Let $X$ be a variety with a regular function $g:X\to{\mathbb{A}}^{1}$. Let $Z:=g^{-1}(0)$, $U=X-Z$, and let $i:Z\hookrightarrow X,j:U\hookrightarrow X$ denote the canonical inclusions. In this setting, one has the following fundamental result due to Beilinson.
Theorem 2.3 (Beilinson).
$\begin{split}B_{g}(U,Z)&=\{(K_{U},K_{Z},c,v)\mid K_{U}\in\operatorname{Perv}(U),K_{Z}\in\operatorname{Perv}(Z)\\ &c:\phi_{g}^{u}K_{U}\to K_{Z},v:K_{Z}\to\phi_{g}^{u}K_{U},cv=T-I\}\end{split}$
In one direction, $K\in\operatorname{Perv}(X)$ goes to $(j^{*}K,\phi_{g}^{u}K,can,var)$. The inverse is more complicated. Beilinson constructs a “maximal extension” functor $\Xi_{g}:\operatorname{Perv}(U)\to\operatorname{Perv}(X)$ which fits into exact sequences
[…]
(2.3.2) $K=H^{0}(i_{*}\psi^{u}_{g}K_{U}\xrightarrow{(c,\alpha)}\Xi_{g}K_{U}\oplus i_{*}K_{Z}\xrightarrow{(v,-\beta)}i_{*}\psi^{u}_{g}K_{U})$
(where this is centered in degree 0).
Remark 2.4.
We observe for later that the first map in (2.3.2) is injective, and the second is surjective.
2.7. Monodromic Sheaves
Let $X$ be a variety over an algebraically closed field $k$ such that $char(k)=0$. Suppose that $X$ comes equipped with a ${\mathbb{G}}_{m}$-action denoted by $\theta:{\mathbb{G}}_{m}\times X\rightarrow X.$ For $\lambda\in k^{\times}$, let $\theta_{\lambda}:=\theta(-,\lambda):X\rightarrow X$, and for $x\in X(k)$, let $w_{x}=\theta(x,-):{\mathbb{G}}_{m}\rightarrow X$.
Example 2.5.
A constructible sheaf (respectively complex of sheaves) ${\mathcal{K}}$ on $X$ is monodromic if $w_{x}^{*}({\mathcal{K}})$ (respectively the cohomology sheaves of $w_{x}^{*}({\mathcal{K}})$) are locally constant for all $x\in X(k)$. We let $Perv_{m}(X)$ denote the full-subcategory of monodromic perverse sheaves.
Lemma 2.6.
([28], 3.2) With notation as above, the following are equivalent:
  • ${\mathcal{K}}$ is monodromic.
Proof.
In the setting of cones and constructible sheaves, this is contained Proposition 3.2 of [28]. The same proof goes through in the current setting. Moreover, the case of complexes can be reduced to sheaves by passing to cohomology sheaves. ∎
Let $X$ be a variety equipped with a ${\mathbb{G}}_{m}$-action, and $h:X\rightarrow{\mathbb{A}}^{1}$ a morphism which is equivariant for the standard ${\mathbb{G}}_{m}$-action on ${\mathbb{A}}^{1}$. Then $Z:=h^{-1}(0)$ and $U$ are invariant under ${\mathbb{G}}_{m}$. In this case, we may consider the categories $\operatorname{Perv}_{m}(X),\operatorname{Perv}_{m}(U),$ and $\operatorname{Perv}_{m}(Z)$.
Theorem 2.7.
With notation as above, $\operatorname{Perv}_{m}(X)$ is equivalent to the full sub-category $B_{g,m}(U,Z)\subset B_{g}(U,Z)$ consisting of objects such that $K_{U}$ and $K_{Z}$ are monodromic.
Proof.
Remark 2.8.
[…]
3. Perverse sheaves on Stacks and Euler characteristics
In this section, we explain the construction of the Euler-Satake charactersitic for constructible sheaves on Deligne-Mumford or DM stacks, and recall some basic facts about perverse sheaves on such stacks. We will usually reserve the symbols ${\mathscr{A}},{\mathscr{M}},{\mathscr{X}},\ldots$ for stacks. A $1$-morphism between stacks will often be referred to simply as a morphism or a map.
@@ 3. Perverse sheaves on Stacks and Euler characteristics
    $k={\mathbb{C}}$, we take constructible sheaf in the analytic topology, and $\Lambda$ is a field.
@@
    $k$ is arbitrary, $char(k)=0$ and we consider constructible sheaves in the lisse-etale topos with $\Lambda={\mathbb{Z}}/\ell^{n}{\mathbb{Z}}$ or ${\mathbb{Q}}_{\ell}$.
In any of the aforementioned settings and when $\Lambda$ is a field, one can define the Euler-Satake characteristic $\chi^{orb}({\mathcal{F}})$ of ${\mathcal{F}}\in{\rm D}^{b}_{c}({\mathscr{X}})$. In the analytic setting and where one considers constructible functions, this was defined by Behrend ([5]). And it was considered by several authors for constant coefficients, starting with Satake in the 1950’s. The Euler-Satake characteristic satisfies the following properties and is uniquely determined by these properties:
@@
    $\chi^{orb}({\mathscr{X}},{\mathcal{F}})\in{\mathbb{Q}}.$
@@
    If $X$ is a scheme, the $\chi^{orb}({\mathscr{X}},{\mathcal{F}})=\chi({\mathscr{X}},{\mathcal{F}})$ i.e. the orbifold Euler characteristic is the usual Euler characteristic. Note: In the analytic setting one should instead take the compactly supported Euler characteristic. These are the same in the algebraic setting.
@@
    Let ${\mathcal{F}}$ be a constructible sheaf on ${\mathscr{X}}$ and ${\mathcal{G}}$ on ${\mathscr{Y}}$. Then $\chi^{orb}({\mathscr{X}}\times{\mathscr{Y}},{\mathcal{F}}\otimes{\mathcal{G}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})\chi^{orb}({\mathscr{Y}},{\mathcal{G}}).$
@@
    Let ${\mathcal{F}}\to{\mathcal{H}}\to{\mathcal{G}}\xrightarrow{[1]}$ be a distinguished triangle in $D_{c}^{b}({\mathscr{X}},\Lambda)$, then $\chi^{orb}({\mathscr{X}},{\mathcal{H}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})+\chi^{orb}({\mathscr{X}},{\mathcal{G}}).$
  7. (vii)
    One has $\chi({\mathscr{X}},{\mathcal{F}}[d])=(-1)^{d}\chi({\mathscr{X}},{\mathcal{F}})$
Proof.
(Proof of existence) Note that the usual (compactly supported) euler characteristic on schemes or complex analytic spaces satisfies these axioms. This extends uniquely to Deligne-Mumford stacks as follows. If ${\mathscr{X}}$ is a DM stack, then generically it is a quotient by a finite etale group scheme. In particular, one has an open ${\mathscr{U}}\subset{\mathscr{X}}$, a scheme $U$ with an action of a finite etale group scheme $G$ such that ${\mathscr{U}}=[U/G]$. On this open ${\mathscr{U}}$, $\chi({\mathscr{U}},{\mathcal{F}}|_{\mathscr{U}})$ is determined by (iii) above. Now one can proceed via induction on dimension and use property (ii). Alternately, one can also stratify ${\mathscr{X}}$ by quotients of the form above. ∎
@@ 3.2. Perverse sheaves on stacks
In ([20]), Laszlo-Olsson define perverse sheaves in the lisse-étale topology of an Artin stack of finite type (over a fixed field $k$ of $char(k)=0$). We refer to loc. cit. for the details of this construction. Below, we recall some basic facts which will be useful in the following. We denote by $\operatorname{Perv}({\mathscr{X}})$ the corresponding category of perverse sheaves (with coefficients in ${\mathbb{Q}}$).
@@ 4. The case of ${}_{N}{\mathscr{A}}_{g}$
Fix $N\geq 3$, and let ${}_{N}{\mathscr{A}}_{g}$ denote the moduli space of principally polarized $g$ dimensional abelian varieties with level $N$ structure over $k$. This is a fine moduli space, and it is nonsingular. Let ${\overline{{}_{N}{\mathscr{A}}_{g}}}$ denote a nonsingular toroidal compactification such that $E={\overline{{}_{N}{\mathscr{A}}_{g}}}-{}_{N}{\mathscr{A}}_{g}$ is a divisor with normal crossings ([10]). In this section, we prove the following theorem:
@@
Let ${\mathcal{P}}$ be a perverse sheaf on $X={}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ where $N_{i}\geq 3$. Then $\chi(X,{\mathcal{P}})\geq 0$
@@
This follows from lemmas 4.3, 4.4 and 4.5 below. ∎
Before proving the theorem, we record the following corollary.
@@
If $X\subseteq{}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ ($N_{i}\geq 3$) is a smooth subvariety, then
$(-1)^{\dim X}\chi(X)\geq 0$
@@
Let $A:={}_{N}{\mathscr{A}}_{g}$ and $\bar{A}={\overline{{}_{N}{\mathscr{A}}_{g}}}$. Then $\Omega_{\bar{A}}^{1}(\log E)$ is nef
@@
Let $f:U\to A$ denote the universal family of abelian varieties, and let $\overline{F}^{1}$ be the associated Hodge bundle over $\bar{A}$. By [10, chap IV, theorem 7.7],
@@
Therefore the lemma follows from theorem 2.2 and [21, thm 6.2.12]. ∎
Lemma 4.4.
Proof.
$\chi(A,P)=\langle\overline{CC(P)},Z\rangle$
where $Z\subset T^{*}_{\bar{A}}(\log E)$ is the zero section. Therefore it is enough to show that this intersection number is nonnegative. Since the log cotangent bundle is nef. Theorem 2.1 yields the desired nonnegativity.
Lemma 4.5.
Let $X_{1}$ and $X_{2}$ be smooth projective varieties with snc divisors $E_{i}$, such that $\Omega_{X_{i}}^{1}(\log E_{i})$ are nef. Then $\Omega_{X_{1}\times X_{2}}^{1}(\log p_{1}^{*}E_{1}+p_{2}^{*}E_{2})$ is nef, where $p_{i}:X_{1}\times X_{2}\to X_{i}$ the projections.
Proof.
This follows from [21, theorem 6.2.12] and the fact that
We conclude the section with some Hodge theoretic variants (which will not be used in the following).
@@
By [13, 0.3A] $X$ is Kähler hyperbolic, therefore the theorem follows from [13, theorem 2.5]. ∎
@@ 5. Euler characteristics of perverse sheaves on ${\mathscr{A}}_{g,n}$
Let $f:(X,E)\to(Y,D)$ be a projective log smooth map of nc log schemes. Set ${\mathcal{X}}=(X,E)$, ${\mathcal{Y}}=(Y,D)$, $\Omega^{1}_{\mathcal{X}}=\Omega_{X}^{1}(\log E)$ etc. We have an exact sequence
@@
This follows from the fact that the nef property is closed under pull backs and extensions [21, thm 6.2.12].
Fix $N\geq 3$. Since ${}_{N}{\mathscr{A}}_{g}$ is a fine moduli space, we have a universal abelian scheme ${}_{N}{\mathscr{A}}_{g,1}\to{}_{N}{\mathscr{A}}_{g}$. Let ${}_{N}{\mathscr{A}}_{n}={}_{N}{\mathscr{A}}_{1}\times_{{}_{N}{\mathscr{A}}_{g}}\ldots\times_{{}_{N}{\mathscr{A}}_{g}}{}_{N}{\mathscr{A}}_{1}$ denote the $n$-fold fibre product. This a fine moduli space, whose $S$-valued points are $g$ dimensional abelian schemes of $S$ with a principal polarization, level $N$ structure, and $n$ sections.
@@
If ${\mathcal{P}}$ be a perverse sheaf on $X={}_{N}{\mathscr{A}}_{g_{1},n_{1}}\times\ldots\times{}_{N}{\mathscr{A}}_{g_{r},n_{r}}$, then $\chi(X,{\mathcal{P}})\geq 0$.
@@
Let
$\overline{X}=\overline{{}_{N}{\mathscr{A}}_{g_{1},n_{1}}}\times\ldots\times\overline{{}_{N}{\mathscr{A}}_{g_{r},n_{r}}}$
be a product of toroidal compactifications. By a theorem of Faltings-Chai [10, chap VI, thm 1.1], we can find an nc log scheme $(P,D)$ with morphism $f:(P,D)\to\overline{X}$ such that
@@
    $P-D=X$ and $f|_{P-D}$ coincides with the product of canonical maps ${}_{N}{\mathscr{A}}_{g_{i},n}\to{}_{N}{\mathscr{A}}_{g_{i}}$.
@@
    $f$ is toroidal [10, chap VI, thm 1.13], and therefore log smooth by [16, prop 3.4].
@@
    $\Omega^{1}_{(P,D)/(\overline{X},E)}$ is the pull back of a Hodge bundle on $\overline{X}$.
The theorem now follows from lemmas 4.4 and 5.1, and theorem 2.2.
@@
Since ${\mathscr{X}}=\prod[{}_{N}{\mathscr{A}}_{g_{i},n_{i}}/PSp_{2g_{i}}({\mathbb{Z}}/N{\mathbb{Z}})]$, this follows from 3.1 (iii) and theorems 4.1 and 4.6.
@@
The theorem cannot be improved to a strict inequality in general, because if $n>0$, then $\chi^{orb}({\mathscr{A}}_{g,n})=0$. However, we do not know if there is an example of $X\subset{\mathscr{A}}_{g}$ with $\chi(X)=0$.
Recall that a morphism $f:X\to Y$ is called semismall if for all $k$
$2k\leq\dim Y-\dim\{y\in Y\mid\dim f^{-1}(y)=k\}$
@@
By [9, prop 4.2.1] $\pi_{*}{\mathbb{Q}}[\dim X]$ is perverse. ∎

### v2

@@ 2.5. Hodge bundles
Suppose that $(Y,D)$ is a projective nc log scheme over $k$ as above. Let $\overline{f}:(X,E)\to(Y,D)$ be a projective log smooth morphism to $(Y,D)$, which restricts to an abelian scheme $f:A\to Y-D$ over YDY-D. Let $\overline{F}^{1}=\overline{f}_{*}\Omega^{1}_{(X,E)/(Y,D)}$.
@@
Without loss of generality, we can assume that $k={\mathbb{C}}$. Switching to the analytic topology, $R^{1}f^{an}_{*}{\mathbb{C}}$ underlies a polarized variation of Hodge structure $(V,F^{\bullet},\ldots)$ of type $\{(1,0),(0,1)\}$, with unipotent monodromy about components of $D$. The Hodge bundle $F^{1}V=f^{an}_{*}\Omega^{1}_{A/Y-D}$ extends to $(\overline{F}^{1})^{an}$ over $Y$. The theorem now follows from the much more general result due to Fujino-Fujisawa-Saito [ffs, theorem 3]. ∎
@@
In this subsection, we explain the construction of the Euler-Satake (or orbifold Euler) characteristic for constructible sheaves on Deligne-Mumford or DM stacks. We will usually reserve the symbols ${\mathscr{A}},{\mathscr{M}},{\mathscr{X}},\ldots$ for stacks. A $1$-morphism between stacks will often be referred to simply as a morphism or a map.
@@
    If $k={\mathbb{C}}$, we take constructible sheaves in the analytic topology, and let $\Lambda$ be a field.
@@
    If $k$ is arbitrary, we consider constructible sheaves in the lisse-etale topos (c.f. [laumon, lo]) with $\Lambda={\mathbb{Z}}/\ell^{n}{\mathbb{Z}}$ or $\mathbb{Q}$$_{\ell}$.
In any of the aforementioned settings and when $\Lambda$ is a field, one can define the Euler-Satake characteristic $\chi^{orb}({\mathcal{F}})$ of ${\mathcal{F}}\in{\rm D}^{b}_{c}({\mathscr{X}},\Lambda)$. In the analytic setting, a version of this, with $\mathcal{F}$} replaced by a constructible function, was defined by Behrend ([Beh]); $\chi^{orb}$ was considered by several authors for constant coefficients, starting with Satake in the 1950’s. The Euler-Satake characteristic is characterized as follows:
Theorem 2.5.
There exists a rational number $\chi^{orb}({\mathscr{X}},{\mathcal{F}})\in{\mathbb{Q}}$ associated to a DM stack $\mathscr{X}$} of finite type over a field $k$ and ${\mathcal{F}}\in{\rm D}^{b}_{c}({\mathscr{X}})$ uniquely determined by the following properties:
@@
    If ${\mathscr{X}}=X$ is a scheme, then $\chi^{orb}({\mathscr{X}},{\mathcal{F}})=\chi(X,{\mathcal{F}})$ i.e. the orbifold Euler characteristic is the usual Euler characteristic.
@@
    Let $\mathcal{F}$} be a constructible sheaf on $\mathscr{X}$} and $\mathcal{G}$} on $\mathscr{Y}$}. Then $\chi^{orb}({\mathscr{X}}\times{\mathscr{Y}},{\mathcal{F}}\boxtimes{\mathcal{G}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})\chi^{orb}({\mathscr{Y}},{\mathcal{G}}).$
@@
    Let ${\mathcal{F}}\to{\mathcal{H}}\to{\mathcal{G}}\xrightarrow{[1]}$ be a distinguished triangle in ${\rm D}_{c}^{b}({\mathscr{X}},\Lambda)$, then $\chi^{orb}({\mathscr{X}},{\mathcal{H}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})+\chi^{orb}({\mathscr{X}},{\mathcal{G}}).$
@@
    One has $\chi^{orb}({\mathscr{X}},{\mathcal{F}}[d])=(-1)^{d}\chi^{orb}({\mathscr{X}},{\mathcal{F}})$
Sketch.
By [laumon, cor 6.1.1] and quasicompactness, we can find a finite open cover $\{{\mathscr{U}}_{1},\ldots,{\mathscr{U}}_{N}\}$, where each ${\mathscr{U}}_{i}\cong[U_{i}/G_{i}]$ is the quotient of a scheme by an etale group scheme. Therefore, we can use induction on $N$ to prove the existence and uniqueness of $\chi^{orb}(X,{\mathcal{F}})$. When $N=1$, axiom (iii) forces $\chi^{orb}([U_{1}/G_{1}],{\mathcal{F}})=\frac{1}{|G_{1}|}\chi(U_{1},{\mathcal{F}}_{1})$, where ${\mathcal{F}}_{1}$ is the pull back of $\mathcal{F}$} to $U_{1}$. Standard properties of Euler characteristics, will show that this number will satisfy the remaining axioms. If $N>1$, set ${\mathscr{Z}}={\mathscr{X}}\setminus{\mathscr{U}}_{1}$, then $\chi^{orb}({\mathscr{Z}},{\mathcal{F}}|_{{\mathscr{Z}}})$ is defined by induction. Axiom (ii) now forces
$\chi^{orb}({\mathscr{X}},{\mathcal{F}})=\chi^{orb}({\mathscr{U}}_{1},{\mathcal{F}}|_{{\mathscr{U}}_{1}})+\chi^{orb}({\mathscr{Z}},{\mathcal{F}}|_{{\mathscr{Z}}})$
Since both terms on the right satisfy the above axioms, the same is true for the left hand side. ∎
@@ 2.7. Perverse sheaves on stacks
In ([lo]), Laszlo-Olsson define perverse sheaves in the lisse-étale topology of an Artin stack of finite type (over a fixed field $k$ of $char(k)=0$). We refer to loc. cit. for the details of this construction. Below, we recall some basic facts which will be useful in the following. We denote by $\operatorname{Perv}({\mathscr{X}})$ the corresponding category of perverse sheaves (with coefficients in $\mathbb{Q}$} or $\mathbb{Q}$$_{\ell}$).
@@ 3. The case of $_{N}{\mathscr{A}}_{g}$
Fix $N\geq 3$, and let $_{N}{\mathscr{A}}_{g}$ denote the moduli space of principally polarized $g$ dimensional abelian varieties with level $N$ structure over $k$. This is a fine moduli space, and it is nonsingular. Let $\overline{{}_{N}{\mathscr{A}}_{g}}$} denote a nonsingular toroidal compactification such that $E={\overline{{}_{N}{\mathscr{A}}_{g}}}-{}_{N}{\mathscr{A}}_{g}$ is a divisor with normal crossings ([fc]). In this section, we prove the following theorem:
@@
Let $\mathcal{P}$} be a perverse sheaf on $X={}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ where $N_{i}\geq 3$. Then $\chi(X,{\mathcal{P}})\geq 0$.
@@
This follows from Lemmas 2.2 and 2.3, and Lemma 3.3 below. ∎
@@
If $X\subseteq{}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ ($N_{i}\geq 3$) is a smooth or lci subvariety, then
$(-1)^{\dim X}\chi(X)\geq 0.$
@@
Let $A:={}_{N}{\mathscr{A}}_{g}$ and $\bar{A}={\overline{{}_{N}{\mathscr{A}}_{g}}}$. Then $\Omega_{\bar{A}}^{1}(\log E)$ is nef.
@@
Let $f:U\to A$ denote the universal family of abelian varieties, and let $\overline{F}^{1}$ be the associated Hodge bundle over $\bar{A}$. By [fc, chap IV, theorem 7.7],
@@
Therefore the lemma follows from theorem 2.4 and [lazarsfeld, thm 6.2.12]. ∎
We conclude the section with some Hodge theoretic variants.
@@
By [gromov, 0.3A] $X$ is Kähler hyperbolic, therefore the theorem follows from [gromov, theorem 2.5]. ∎
@@ 4. Euler characteristics of perverse sheaves on ${\mathscr{A}}_{g,n}$
Let $f:(\overline{X},E)\to(\overline{Y},D)$ be a projective log smooth map of nc log schemes. Set ${\mathcal{X}}=(\overline{X},E)$, ${\mathcal{Y}}=(\overline{Y},D)$, $\Omega^{1}_{\mathcal{X}}=\Omega_{\overline{X}}^{1}(\log E)$ etc. We have an exact sequence
@@
This follows from the fact that the nef property is closed under pull backs and extensions [lazarsfeld, thm 6.2.12].
Fix $N\geq 3$. Since $_{N}{\mathscr{A}}_{g}$ is a fine moduli space, we have a universal abelian scheme $_{N}{\mathscr{A}}_{g,1}\to{}_{N}{\mathscr{A}}_{g}$. Let $_{N}{\mathscr{A}}_{g,n}={}_{N}{\mathscr{A}}_{g,1}\times_{{}_{N}{\mathscr{A}}_{g}}\ldots\times_{{}_{N}{\mathscr{A}}_{g}}{}_{N}{\mathscr{A}}_{g,1}$ denote the $n$-fold fibre product. This a fine moduli space, whose $S$-valued points are $g$ dimensional abelian schemes of $S$ with a principal polarization, level $N$ structure, and $n$ sections. We let ${\mathscr{A}}_{g}=[{}_{N}{\mathscr{A}}_{g}/PSp_{2g}({\mathbb{Z}}/N{\mathbb{Z}})]$, ${\mathscr{A}}_{g,n}=[{}_{N}{\mathscr{A}}_{g,n}/PSp_{2}({\mathbb{Z}}/N{\mathbb{Z}})]$ denote the moduli stacks of principally polarized $g$ dimensional abelian schemes (with $n$ sections).
@@
Let $X={}_{N}{\mathscr{A}}_{g_{1},n_{1}}\times\ldots\times{}_{N}{\mathscr{A}}_{g_{r},n_{r}}$. Then there is an nc log scheme $(\overline{X},E)$ such that $\Omega_{(\overline{X},E)}^{1}$ is nef, $\overline{X}$ is proper and $X=\overline{X}\setminus E$. If $\mathcal{P}$} is a perverse sheaf on $X$, then $\chi(X,{\mathcal{P}})\geq 0$.
@@
By work of Mumford et. al., we can find a smooth toroidal compactification $\overline{{}_{N}{\mathscr{A}}_{g_{i}}}$ of $_{N}{\mathscr{A}}_{g_{i}}$ such that boundary has simple normal crossings [fc, chap IV]. Let
$\overline{Y}=\overline{{}_{N}{\mathscr{A}}_{g_{1}}}\times\ldots\times\overline{{}_{N}{\mathscr{A}}_{g_{r}}}$
be a product of these compactifications, and $D$ denote the simple normal crossings boundary divisor. By a theorem of Faltings-Chai [fc, chap VI, thm 1.1], we can find an nc log scheme $(\overline{X},E)$ with a morphism $f:(\overline{X},E)\to(\overline{Y},D)$ such that
@@
    $\overline{X}-E=X$ and $f|_{X}$ coincides with the product of canonical maps $_{N}{\mathscr{A}}_{g_{i},n}\to{}_{N}{\mathscr{A}}_{g_{i}}$.
@@
    $f$ is toroidal [fc, chap VI, thm 1.13], and therefore log smooth by [kato, prop 3.4].
@@
    $\Omega^{1}_{(\overline{X},E)/(\overline{Y},D)}$ is the pull back of a Hodge bundle on $\overline{Y}$.
The theorem now follows from Lemmas 2.2, 3.3 and 4.1, and Theorem 2.4.
@@
Since ${\mathscr{X}}=\prod[{}_{N}{\mathscr{A}}_{g_{i},n_{i}}/PSp_{2g_{i}}({\mathbb{Z}}/N{\mathbb{Z}})]$, this follows from Theorem 2.5 (iii) and Theorems 3.1, 3.4 and 4.2.
@@
The theorem cannot be improved to a strict inequality in general, because if $n>0$, then $\chi^{orb}({\mathscr{A}}_{g,n})=0$ (since $_{N}{\mathscr{A}}_{g,n}$ is an abelian scheme). However, we do not know if there is an example of $X\subset{\mathscr{A}}_{g}$ with $\chi(X)=0$.
Recall that a morphism $f:X\to Y$ is called semismall if for all $m$
$2m\leq\dim Y-\dim\{y\in Y\mid\dim f^{-1}(y)=m\}$
@@
By [dm, prop 4.2.1] $\pi_{*}{\mathbb{Q}}[\dim X]$ is perverse. ∎